CLIMMRC(5) CLIMMRC(5) ˜ /.climm/climmrc−climm climm.[],.,˜ /.climm/. CONNECTIONS Connection()ICQ-.:ICQ""--.:FIFO. type[auto] ,:server()ICQ,peer()--,remote()FIFO.auto(),. UIN-;UIN-. version .5,78--6,78.FIFO-. server .login.icq.com.--.fifo-FIFO. port,""--. uinUIN UIN-UIN.--FIFO. password[] .--.,.:BeOS. status .: 0 1 5 17 19 32 256 --,: 1.. 2,SOCKS. 4.a. FIFO-. GENERAL General()SOCKS5,. s5use 01SOCKS5. s5host socks5. s5port socks5"". s5auth 01. climm 1 CLIMMRC(5) CLIMMRC(5) s5name socks5. s5pass socks5. climmrc: encodingfile .climm(utf8),,.,climm. encodinglocal .. encodingremote -,,.,. format .,1. : verbose . autoaway "".0. Screenwidth .,0,80.. setdelbson|off on(),deletebackspace.off(),delete. setfunnyon|off

"".

setcoloron|off . sethermiton|off on(),-. setlogon|off . setlogonoffon|off

,""/"".

setautoon|off (reply). setuinprompton|off :. setautosaveon|off climmrc. setautofingeron|off UIN-. setlinebreak ,:simple():,break():,indent():smart():. settabssimple|cycle|cycleall .climm-simple(),msg,..cycle()e,-.,msg.cycleall() cycle(),,. climm 2 CLIMMRC(5) CLIMMRC(5) setsilent ,on(),complete(). chat.-1,49climm(). ,: colorscheme . color .:none(),server(),client(),message(),contact(),sent(),ack(),error(), debug()incoming(),:black(),red(),green(),yellow(),blue(),magenta(), cyan(-),white(),none(),bold()(bold,,),. logplace| ,-,-.,,/. soundon|beep|off|event "".on()beep(),off(),event()"". event.: 1.IM,icq. 2.UIN,0. 3.,. 4.global.5.,msg(),on(),off(),beep()status(),on()off().. 6.,,,,,0. 7.. . auto.away(),na(),dnd(),occ(),inv(),ffc(). STRINGS Strings. alter<><> ..climmcmds(7).. :. alias<><> .%s,,.0.4.10. GROUP Group()-.: server<><uin> -,icqv88ICQicqv55ICQ,UINuin.o,. label<> -.--uin,-. id<>.0.-. entry<><uin> UIN-uin-. CONTACTS Contacts()()-.. [*][˜ ][ˆ ]uin UIN-uin.*.˜ .ˆ .uin,. climm(1),climmcmds(7) climm 3 CLIMMRC(5) CLIMMRC(5) (JamesMorrison)<ja2morrison@student.math.uwaterloo.ca>.,,(Ru¨ digerKuhlmann) <climm@ruediger-kuhlmann.de>.:<lecicct@ptt.yu>. climm 4